An educational psychologist wishes to know the mean number of words a third grader can read per minute. She wants to ensure that the estimate has an error of at most 0.24 words per minute. A previous study found that the mean was 29.6 words per minute. Assuming that the variance is 29.16, what is the minimum number of third graders that must be included in a sample to construct the 90% confidence interval? Round your answer up to the next integer.
Answer #1

The sample size formula is given as below:

n = (Z*σ/E)^2

We are given

σ^2 = 29.16

σ = sqrt(29.16) = 5.4

Confidence level = 90%

Critical Z value = 1.6449

(by using z-table)

Margin of error = E = 0.24

The sample size is given as below:

n = (Z*σ/E)^2

n = (1.6449*5.4/0.24)^2

n = 1369.759

Required sample size = 1370